OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form.

OpenStudy (anonymous):

y-intec form is y=mx+b first multiply -2 by x and by -5 u will get y-4=-2x+10 +4 +4 y=-2x+14